# Machine learning for advancing low-temperature plasma modeling and simulation

### v1

@@ 3.2 Plasma Physics
Describing the evolution of dynamical systems (or sub-systems) accurately is an omnipresent task in scientific computing. The time integration schemes of differential operators have been correspondingly considered in different fields (e.g., pairing Runge–Kutta integration schemes with ANNs [181, 55, 157, 182]). For the time integration of multiscale physical systems, Liu et al.[157] have suggested an hierarchical time-steppers (HiTSs) framework. The approach builds on the decomposition of the system dynamics into a hierarchy of prediction steps that can be arranged and interpolated in time. It has been shown to outperform other methods based on LSTM, ESN, and clockwork RNNs for the Lorenz system, 2D fluid flow past a cylinder, the Kuramoto–Sivashinsky (KS) equation, as well as music and video sequences. An illustrative example comparison of the prediction output is given in Figure 5. LTP modeling and simulation is facing essentially similar problems, suggesting it may similarly benefit from application of the HiTSs method (and also the other referenced methods) in the description of the multiscale dynamics.
An approach using the concept of dynamic mode decomposition (DMD) to obtain low-dimensional linear dynamics of given data has been applied to magnetic confinement fusion plasma simulations by De Pascuale et al.[183] The obtained dynamic modes are evolved individually solving a linear ODE. The constructed linear time advance operator based on DMD has been successfully applied for the analysis of several 1D diffusion transport problems, as well as for dynamics of 1D profiles at the divertor targets of a tokamak plasma. This scheme could be similarly applied to complex LTP processes.
The solution to Newton’s equations for particle based simulations has been explored by Kadupitiya et al. using LSTM networks.[184] The accurate applicability has been investigated for few-particle systems subject to various interaction potentials (e.g., Lennard–Jones). Energy-conserving dynamics up to 4000 larger timesteps compared to a Verlet integration have been demonstrated. Following the concept of graph neural networks,[98, 99] which embed the interaction dynamics of nodes (e.g., particles) in their learnable edge weights (e.g., interaction), the integration of ODEs in an Hamiltonian framework, have been discussed by Sanchez-Gonzalez et al.[185, 186] The proposed Hamiltonian ODE graph network trained on Runge–Kutta scheme data generalizes the integrators to accurately represent the Hamiltonian dynamics. In the context of many-particle simulations, Li et al.[187] have proposed a GNN accelerated MD framework that predicts the interaction forces for the system, given its momentary state (e.g., atom positions and types). It has been argued that the scheme can be easily integrated into existing numerical integrators. Due to the consideration of multi-particle interactions, this scheme may be specifically relevant in PIC simulations should Coulomb interactions be important.
@@
Concerning data-driven process control and MPC it should be noted that the physics-informed VM methodology for LTP etching processes previously introduced may be adapted based on physical or data-driven model data.[19, 190, 20] Despite a possible discrepancy between models and reality, in particular systematic deviations may be straightforwardly taken into account in a data-driven approach. A learnable translation layer between these may suffice to capture and combine the relevant information by means of data-fusion. This may allow data-driven MPC with limited observables (cf. also PC-EKF[148, 149]). Additionally, MPC schemes based on LSTM networks that have been successfully demonstrated in magnetic confinement fusion plasmas subject to inherent complex nonlinear dynamics by Degrave et al.[191] may be promising candidates for adaptation in LTP processing.
@@ 3.3 Plasma chemistry
The adoption of detailed mechanisms for chemical kinetics often poses two types of challenges. First, the number of degrees of freedom is large; and second, the dynamics is characterized by a wide range of time scales. This has motivated the development of several techniques for reducing the complexity of such kinetic models where only a few variables are considered in the simplified model. In [192], Grassi and co-authors have used AEs to reduce complex chemical data sets and decrease the computational time associated with numerical solutions of ODEs for a large number of species and reactions. In particular, an encoder is used to reduce the dimensionality of the chemical state space into a latent (lower-dimensional) space. Analogously to the chemical network in the original space, that is represented by the system of ODEs, the chemical network in the latent space can be evolved in time with a different set of ODEs which has got a significantly smaller number of dimensions. Hence, a decoder is used to transform the variable from the latent space back to the original space. In [192], the model enables a compression of chemical networks composed of 224 reactions and 29 species into a network with 12 reactions and 5 species that are evolved in time using a standard ODE solver. Their work has demonstrated that a 65-times speed-up can be achieved. Tang and Turk [193] have proposed a combination of AE and neural ODE to model the temporal evolution of chemical kinetics in a reduced subspace. The architecture of the network is similar to the one by Grassi and co-authors [192]. In their work, a 10-fold speed-up compared to commonly used astro-chemistry solver for a 9-species primordial network has been achieved, while maintaining accuracy across a wide range of density and temperature. As described in subsection 2.2, it is well known that, due to the presence of fast and slow dynamics, the chemical systems are characterized by low-dimensional manifolds in the concentration space. Chiavazzo and co-authors[194] applied the Diffusion Map (DMAP) approach for constructing reduced kinetics models for combustion applications. The DMAP approach can be considered as a non-linear counterpart of PCA that can be used for searching a low-dimensional embedding of a high-dimensional chemical data set.
@@
Stiff equations for chemical kinetics are computationally expensive to solve numerically. The efficacy of ML for applications that need to solve systems of ODEs or PDEs has been explored in many areas, such as computational fluid dynamics (CFD) and combustion. Goswami and co-authors[195] have developed a neural operator-based surrogate model (DeepONet) to efficiently solve stiff chemical kinetics. The model was applied to a chemical reaction system for CO/H burning of syngas, which contains 11 species and 21 reactions, and a temporally developing planar CO/H jet flame (turbulent flame) using the same syngas mechanism. It has been shown that the DeepONet, once trained, can accurately integrate the thermochemical state for arbitrarily large time advancements, leading to significant computational gains compared to stiff integration schemes. Moreover, both results obtained with DeepONet and AE-based DeepONet are computationally very efficient compared to conventional CFD solvers. One of the advantages of this approach is the possibility to be included in a CFD solver for solving such stiff chemically reacting problems, which can drastically reduce the computational cost. Campoli and co-authors[196] have performed state-to-state numerical simulations of high speed reacting gas flow using data-driven ML regression models. The models were applied to a system of equations for a 1D reacting flow of a five-component air mixture including a total of 122 excited states of N, O, NO, N, and O species with detailed vibrational kinetics. Several methods have been compared for estimation of the state-to-state relaxation rate terms, such as Kernel Ridge (KR), Support Vector Machines (SVM), k-Nearest Neighbor, Gaussian Processes, Decision Tree (DT), Random Forest, Extremely Randomized Trees, Gradient Boosting, Histogram-Based Gradient Boosting, and MLP. From their comparison, it has been found that KR reports the lowest error levels while the SVM reports the highest. Comparable error levels are reported by the remaining algorithms; however, there are noticeable differences in the prediction time, where DT is found to be the fastest algorithm. Finally, the solution of the state-to-state Euler system of equations was inferred by means of a deep neural network by-passing the use of the solver. As a result of this study, it has been been demonstrated that embedding of ML algorithms into ODE solvers offers a speed-up of several orders of magnitude. Chemical Reactor Networks (CRN) have been investigated by Savarese and co-authors [197] for performing faster CFD simulations of combustion systems. In their work, a novel automatic data-driven method for the design of CRN models has been proposed. The method is based on a combination of unsupervised clustering and graph scanning algorithms. Results of their work show that the novel methodology is capable of extracting equivalent CRN models of realistic combustion devices from CFD data. Furthermore, the unsupervised clustering algorithm was able to automatically detect important regions within the combustor domain, such as inlets, inflame and post-flame reactors. Overall, the approach is promising as it has the potential to drive towards the development of ROMs without a huge amount of user-based knowledge. In [198], Mao and co-authors have developed a novel CFD solver that combines the multi-functionalities of the OpenFOAM library[199] with the ML framework Torch[200] and chemical kinetics program Cantera[201]. In their solver, DNN are used to accelerate chemistry solvers and improve the simulation efficiency. As a result, a speed-up of two orders of magnitude is achieved in a simple hydrogen ignition case when performed on a medium-end GPU. This work highlights the potential of the integration of ML techniques within current libraries and computing architecture for detailed CFD simulations of complex hydrocarbon fuels.
@@ 3.4 Plasma-surface interaction
However, in the following a joint, overarching challenge relevant to all fields of LTP processing will be discussed.
i) Scale-bridging
@@
The intrinsic time and length scales of the two states of matter (i.e., plasma, solid-state) involved differ in orders of magnitude. From the material scientist point of view, who resolves atomic time and length scales, PSIs have to be perceived as individual or a sequence of impinging particles, neglecting the true spatiotemporal correlations. Multiscale modeling approaches which repeatedly substitute short time and length scales are, however, always accompanied with an information loss of the fundamental atomic phenomena. ML is assumed to provide an alternative.
@@
Information on the electron dynamics, specifically, on the secondary electron emissions are essential for LTP modeling and simulation. Though experimental as well as theoretical approaches have been pursued in the past, the availability of such coefficients is scarce. More recent efforts include the combination of both (i.e., PIC/MCC simulations, phase resolved OES) and was framed as -CAST.[204, 205] Moreover, a review of Monte Carlo methods to simulate the electron scattering in solids was put together by Chang et al.,[206] who studied the low-energy electron deposition in W surfaces. Electron force field simulations have been demonstrated to predict electron emissions due to the Auger process.[207]. Auger neutralization was also described by Pamperin et al.,[208] who have formulated a generic quantum-kinetic approach, that is based on Anderson-Newns-type effective Hamiltonians. Bronold and Fehske[209] have derived an equation for the electron emission yield in case of low energy electrons impinging onto metal surfaces. An invariant embedding principle was employed. Such approaches could complement each other when combined by a generalizing AI.
@@
Information on the atomic scale have been limited for a long time by a compromise of accuracy and computational efficiency, effectively restricting the physical fidelity of related case studies. Recent advances by means of ML interaction potentials have indicated to eventually overcome this dilemma. However, only a small number of such works have been published so far in the context of LTP modeling and simulation (cf. subsection 2.3). It is therefore essential to further encourage their usage to get used them, enabling more likely developments that meet demands specific to plasma processing. The burden of setting up suitable ML interaction potentials could be bypassed at least to some extent by joining available ones with more traditional methods. For example, the Ziegler-Biersack-Littmark potential[210] and all-electron DFT repulsive potential, DMol, have been merged with ML interaction potentials (e.g., GAP), which were trained to equilibrium properties, to account for screened nuclear repulsion during high-energy collisions when radiating Si or Al.[211, 115, 212, 110] Caro et al.[213, 214] conducted MD simulations with the GAP to study the ion beam deposition of C thin films. Studying a complementary sputter deposition can be readily achieved by introducing Ar ions to the process, whose interactions are sufficiently well described by traditional means. An overview of and introduction to ML interaction potentials for the popular Large-scale Atomic/Molecular Massively Parallel Simulator (LAMMPS) has been presented by Thompson et al.[215]
@@
The increased efficiency of such approaches allows indeed for a more thorough but still accurate parameter space exploration, which is of major importance since most PSIs (e.g., surface chemical reactions, ion radiation) have got too large degrees of freedom (e.g., reaction path ways, 0D to 3D defect structures) to straightforwardly study. Remedies combine ML algorithms for clustering (e.g., Bayesian Gaussian Mixture model), dimensionality reduction (e.g., harmonic linear discriminant analysis), or surrogate modeling (e.g., PSNN) with high-throughput simulations (e.g., randomized trajectories,[136] amorphous surface site melt-quenching[109]) or enhanced sampling methods (e.g., variationally enhanced sampling, metadynamics).[216, 104, 109, 136]. But GNN have been demonstrated to significantly accelerate atomic simulations too.[97] An active learning scheme as applied by Diaw et al.[174] when conducting MD simulations of warm dense matter to setup a corresponding surrogate model may be considered to secure spending available resources most effectively.
@@
Whereas samples in the parameter space are naturally related with different process conditions, they may also be viewed as instances or states of a system (surface patch), which it may they evolve through during operation. Lim et al.[216] formulated an according transition state modeling approach[217, 218, 219] for the long-time scale (i.e., of the order of microseconds) restructuring of Pd deposited on Ag. States were found by detecting relevant positional changes of adatoms during MD simulations, clustering them (unsupervised), and eventually sorting them (supervised). Gergs et al.[136] used a PSNN to generalize on previously sampled but disjoined system state trajectories. The system alternately underwent PSIs and diffusion processes for given experimental discharge conditions, continuously evolving the system state (e.g., composition, point defect populations) for process times of up to 45 minutes in merely 34 GPU hours. Following a similar line of thought, such instances or states could also be understood a surface patches, which may be assembled to setup a PSI ML surrogate model on the feature scale. Kim et al.[128] suggest a similar usage for etching high-aspect ratio feature patterns. However, such an approach may also be of significant interest for research on plasma catalysis, taking complex large-scale surfaces into account. Joining such ML models, which are created in a bottom-up style (starting with simulations on the microscopic scale), with a complementary top-down approach (starting with experiments on the macroscopic scale) may ultimately enable a fidelity beyond DFT.

### v2

@@ 3.2 Plasma Physics
Describing the evolution of dynamical systems (or sub-systems) accurately is an omnipresent task in scientific computing. The time integration schemes of differential operators have been correspondingly considered in different fields (e.g., pairing Runge–Kutta integration schemes with ANNs [184, 55, 159, 185]). For the time integration of multiscale physical systems, Liu et al.[159] have suggested an hierarchical time-steppers (HiTSs) framework. The approach builds on the decomposition of the system dynamics into a hierarchy of prediction steps that can be arranged and interpolated in time. It has been shown to outperform other methods based on LSTM, ESN, and clockwork RNNs for the Lorenz system, 2D fluid flow past a cylinder, the Kuramoto–Sivashinsky (KS) equation, as well as music and video sequences. An illustrative example comparison of the prediction output is given in Figure 5. LTP modeling and simulation is facing essentially alike problems, suggesting it may similarly benefit from application of the HiTSs method (and/or other methods referenced) in the description of their multiscale dynamics.
An approach using the concept of dynamic mode decomposition (DMD) to obtain low-dimensional linear dynamics of given data has been applied to magnetic confinement fusion plasma simulations by De Pascuale et al.[186] The obtained dynamic modes are evolved individually solving a linear ODE. The constructed linear time advance operator based on DMD has been successfully applied for the analysis of several 1D diffusion transport problems, as well as for the dynamics of 1D profiles at the divertor targets of a tokamak plasma. This scheme could be similarly applied to complex LTP processes.
The solution to Newton’s equations for particle based simulations has been explored by Kadupitiya et al. using LSTM networks.[187] The accurate applicability has been investigated for few-particle systems subject to various interaction potentials (e.g., Lennard–Jones). Energy-conserving dynamics up to 4000 larger timesteps compared to a Verlet integration have been demonstrated. Following the concept of GNNs,[100, 101] which embed the interaction dynamics of nodes (e.g., particles) in their learnable edge weights (e.g., interaction), the integration of ODEs in an Hamiltonian framework, have been discussed by Sanchez-Gonzalez et al.[188, 189] The proposed Hamiltonian ODE graph network trained on Runge–Kutta scheme data generalizes the integrators to accurately represent the Hamiltonian dynamics. In the context of many-particle simulations, Li et al.[190] have proposed a GNN accelerated MD framework that predicts the interaction forces for the system, given its momentary state (e.g., atom positions and types). It has been argued that the scheme can be easily integrated into existing numerical integrators. Due to the consideration of multi-particle interactions, this scheme may be specifically relevant in PIC simulations should Coulomb interactions be important.
@@
Concerning data-driven process control and MPC it should be noted that the plasma information based VM methodology for LTP etching processes previously introduced may be adapted based on physical or data-driven model data.[19, 193, 20] Despite a possible discrepancy between models and reality, in particular systematic deviations may be straightforwardly taken into account in a data-driven approach by means of discrepancy learning. A learnable translation layer may suffice to capture and combine the relevant information by means of data-fusion. This may enable data-driven MPC with limited observables (cf. also PC-EKF[150, 151]). Additionally, MPC schemes based on LSTM networks that have been successfully demonstrated in magnetic confinement fusion plasmas subject to inherent complex nonlinear dynamics by Degrave et al.[194] may be promising candidates for adaptation in LTP processing.
@@ 3.3 Plasma chemistry
The adoption of detailed mechanisms for chemical kinetics often poses two types of challenges. First, the number of degrees of freedom is large; and second, the dynamics is characterized by a wide range of time scales. This has motivated the development of several techniques for reducing the complexity of such kinetic models where only a few variables are considered in the simplified model. Grassi and co-authors[195] have used AEs to reduce complex chemical data sets and decrease the computational time associated with numerical solutions of ODEs for a large number of species and reactions. In particular, an encoder is used to reduce the dimensionality of the chemical state space into a latent (lower-dimensional) space. Analogously to the chemical network in the original space, that is represented by the system of ODEs, the chemical network in the latent space can be evolved in time with a different set of ODEs which has got a significantly smaller number of dimensions. Hence, a decoder is used to transform the variable from the latent space back to the original space. The proposed model enables a compression of chemical networks composed of 224 reactions and 29 species into a network with 12 reactions and 5 species that are evolved in time using a standard ODE solver. Their work has demonstrated that a 65-times speed-up can be achieved. Tang and Turk [196] have proposed a combination of AE and neural ODE to model the temporal evolution of chemical kinetics in a reduced subspace. The architecture of the network is similar to the one by Grassi and co-authors [195]. In their work, a 10-fold speed-up compared to commonly used astro-chemistry solver for a 9-species primordial network has been achieved, while maintaining accuracy across a wide range of density and temperature. As described in subsection 2.2, it is well known that, due to the presence of fast and slow dynamics, the chemical systems are characterized by low-dimensional manifolds in the concentration space. Chiavazzo and co-authors[197] applied the Diffusion Map (DMAP) approach for constructing reduced kinetics models for combustion applications. The DMAP approach can be considered as a non-linear counterpart of PCA that can be used for searching a low-dimensional embedding of a high-dimensional chemical data set.
@@
Stiff equations for chemical kinetics are computationally expensive to solve numerically. The efficacy of ML for applications that need to solve systems of ODEs or PDEs has been explored in many areas, such as computational fluid dynamics (CFD) and combustion. Goswami and co-authors[198] have developed a neural operator-based surrogate model (DeepONet) to efficiently solve stiff chemical kinetics. The model was applied to a chemical reaction system for CO/H burning of syngas, which contains 11 species and 21 reactions, and a temporally developing planar CO/H jet flame (turbulent flame) using the same syngas mechanism. It has been shown that the DeepONet, once trained, can accurately integrate the thermochemical state for arbitrarily large time advancements, leading to significant computational gains compared to stiff integration schemes. Moreover, both results obtained with DeepONet and AE-based DeepONet are computationally very efficient compared to conventional CFD solvers. One of the advantages of this approach is the possibility to be included in a CFD solver for solving such stiff chemically reacting problems, which can drastically reduce the computational cost. Campoli and co-authors[199] have performed state-to-state numerical simulations of high speed reacting gas flow using data-driven ML regression models. The models were applied to a system of equations for a 1D reacting flow of a five-component air mixture including a total of 122 excited states of N, O, NO, N, and O species with detailed vibrational kinetics. Several methods have been compared for estimation of the state-to-state relaxation rate terms, such as kernel ridge (KR), SVM, k-nearest neighbor, Gaussian processes, decision tree (DT), random forest, extremely randomized trees, gradient boosting, histogram-based gradient boosting, and MLP. From their comparison, it has been found that KR reports the lowest error levels while the SVM reports the highest. Comparable error levels are reported by the remaining algorithms; however, there are noticeable differences in the prediction time, where DT is found to be the fastest algorithm. Finally, the solution of the state-to-state Euler system of equations was inferred by means of a deep neural network by-passing the use of the solver. As a result of this study, it has been been demonstrated that embedding of ML algorithms into ODE solvers offers a speed-up of several orders of magnitude. Chemical reactor networks (CRNs) have been investigated by Savarese and co-authors [200] for performing faster CFD simulations of combustion systems. In their work, a novel automatic data-driven method for the design of CRN models has been proposed. The method is based on a combination of unsupervised clustering and graph scanning algorithms. Results of their work show that the novel methodology is capable of extracting equivalent CRN models of realistic combustion devices from CFD data. Furthermore, the unsupervised clustering algorithm was able to automatically detect important regions within the combustor domain, such as inlets, inflame and post-flame reactors. Overall, the approach is promising as it has the potential to drive towards the development of ROMs without a huge amount of user-based knowledge. Mao and co-authors[201] have developed a novel CFD solver that combines the multi-functionalities of the OpenFOAM library[202] with the ML framework Torch[203] and chemical kinetics program Cantera[204]. In their solver, DNNs are used to accelerate chemistry solvers and improve the simulation efficiency. As a result, a speed-up of two orders of magnitude is achieved in a simple hydrogen ignition case when performed on a medium-end GPU. This work highlights the potential of the integration of ML techniques within current libraries and computing architecture for detailed CFD simulations of complex hydrocarbon fuels.
@@ 3.4 Plasma-surface interaction
In the following, overarching challenges relevant to all fields of LTP processing are discussed. The intrinsic time and length scales of the two states of matter (i.e., plasma, solid-state) differ in orders of magnitude. From a material scientist point of view, who resolves atomic time and length scales, PSIs have to be perceived as individual or a sequence of impinging particles, neglecting the true spatiotemporal correlations. Multiscale modeling approaches which repeatedly substitute short time and length scales are, however, always accompanied with an information loss of the fundamental atomic phenomena. ML is assumed to provide an alternative.
@@
For electrons, information on their dynamics, specifically, on the secondary electron emissions are essential for LTP modeling and simulation.[207] Although experimental as well as theoretical approaches have been pursued in the past, the availability of such coefficients is scarce. More recent efforts include the combination of both (i.e., PIC/MCC simulations, phase resolved OES) and was framed as -CAST.[208, 209] Moreover, a review of Monte Carlo methods to simulate the electron scattering in solids was put together by Chang et al.,[210] who studied the low-energy electron deposition in W surfaces. Electron force field simulations have been demonstrated to predict electron emissions due to the Auger process.[211]. Auger neutralization was also described by Pamperin et al.,[212] who have formulated a generic quantum-kinetic approach, that is based on Anderson-Newns-type effective Hamiltonians. Bronold and Fehske[213] have derived an equation for the electron emission yield in case of low energy electrons impinging onto metal surfaces. An invariant embedding principle was employed. It is argued that such approaches or combinations thereof could synergistically complement each other, in particular when paired with corresponding ML methods.
@@
Information on the atomic scale have been limited for a long time by a compromise of accuracy and computational efficiency, effectively restricting the physical fidelity of related case studies. Recent advances by means of ML interaction potentials have indicated to eventually overcome this dilemma. However, only a small number of such works have been published so far in the context of LTP modeling and simulation (cf. subsection 2.3). It is therefore essential to further encourage their usage, enabling more likely developments that meet demands specific to plasma processing. The burden of setting up suitable ML interaction potentials could be bypassed at least to some extent by joining available ones with more traditional methods. For example, the Ziegler–Biersack–Littmark potential[214] and all-electron DFT repulsive potential, DMol, have been merged with ML interaction potentials (e.g., GAP), which were trained to equilibrium properties, to account for screened nuclear repulsion during high-energy collisions when radiating Si or Al.[215, 117, 216, 112] Caro et al.[217, 218] conducted MD simulations with the GAP to study the ion beam deposition of C thin films. Studying a complementary sputter deposition can be readily achieved by introducing Ar ions to the process, whose interactions are sufficiently well described by traditional means. An overview of and introduction to ML interaction potentials for the popular Large-scale Atomic/Molecular Massively Parallel Simulator (LAMMPS) has been presented by Thompson et al.[219]
@@
The increased efficiency of such approaches allows for a more thorough but still accurate parameter space exploration. This is of major importance since most PSIs (e.g., surface chemical reactions, ion radiation) are subject to excessive degrees of freedom and cannot straightforwardly be evaluated (e.g., reaction path ways, 0D to 3D defect structures). Remedies combine ML algorithms for clustering (e.g., Bayesian Gaussian mixture model), dimensionality reduction (e.g., harmonic linear discriminant analysis), or surrogate modeling (e.g., PSNN) with high-throughput simulations (e.g., randomized trajectories,[138] amorphous surface site melt-quenching[111]) or enhanced sampling methods (e.g., variationally enhanced sampling, metadynamics).[220, 106, 111, 138]. Notably, GNN have also been demonstrated to significantly accelerate atomic simulations.[99] An active learning scheme as applied by Diaw et al.[177] when conducting MD simulations of warm dense matter to setup a corresponding surrogate model may be considered to secure spending available resources most effectively.
@@
Whereas samples in the parameter space are naturally related to different process conditions, they may also be viewed as instances or states of a system (surface patch), which may evolve during operation. Lim et al.[220] formulated an according transition state modeling approach[221, 222, 223] for the long-time scale (i.e., of the order of microseconds) restructuring of Pd deposited on Ag. States were found by detecting relevant positional changes of adatoms during MD simulations, clustering them (unsupervised), and eventually sorting them (supervised). Gergs et al.[138] used a PSNN to generalize on previously sampled but disjoined system state trajectories. The system alternately underwent PSIs and diffusion processes for given experimental discharge conditions, continuously evolving the system state (e.g., composition, point defect populations). Following a similar line of thought, single instances or states could also be understood as surface patches, which may be assembled to setup a PSI ML surrogate model on the feature scale. Kim et al.[130] suggest a similar usage for etching high-aspect ratio feature patterns. However, such an approach may also be of significant interest for research on plasma catalysis, taking complex large-scale surfaces into account. Joining such ML models, which are created in a bottom-up style (starting with simulations on the microscopic scale), with a complementary top-down approach (starting with experiments on the macroscopic scale) may ultimately enable a fidelity beyond DFT.